# Relative Overfitting and Accept-Reject Framework

### v1

@@ B.1 Larger models have more optional tokens
The number of tokens that the model may generate across all datasets increases with the increase in model parameters.
B.2 Larger models have more optional tokens
@@
From Figure 4, it can be observed that under the condition of the same main $\theta_{LLM}$ parameters, the optimal weight ratio $\alpha$ for each model shows an increasing trend with the rise in the parameter ratio.
@@ C Experiments in the rest of the machine learning field
In this appendix, we detail the experimental setups employed. For image recognition, a suite of pre-trained ResNet models of varying sizes (ResNet-18 to ResNet-152) 64 was utilized. These models were subsequently fine-tuned for one epoch on the CIFAR-10, CIFAR-10067 , and Tiny ImageNet-200 datasets 59 ; 60 ; 61 ; 62 ; 63 to assess classification efficacy across varying complexity levels. Concurrently, for the protein sequence prediction task, the ESM-265 model was employed. This task focuses on predicting protein sequences, which holds significant societal importance, including applications such as the design of novel protein sequences for therapeutic or industrial purposes. The model’s predictive capabilities were benchmarked on the UniRef5066 dataset.
The experimental results demonstrate that the ideas and framework we have proposed can be universally applied to other areas of machine learning.
@@ C.1 Computer Vision
In our experiments, we employed a consistent post-training approach across CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. We fine-tuned ResNet series models (from ResNet18 to ResNet152) pre-trained on ImageNet. The fine-tuning process began by replacing the final fully connected layer, adjusting the output dimensions to match the respective datasets: 10 classes for CIFAR-10, 100 for CIFAR-100, and 200 for Tiny-ImageNet. All images were resized to 224×224 pixels and normalized using ImageNet parameters (mean=[0.485, 0.456, 0.406], std=[0.229, 0.224, 0.225]). We utilized SGD optimizer with a learning rate of 0.001 and momentum of 0.9, typically running for one epoch. Performance was evaluated using Top-1 accuracy metrics for both individual models and ensembles.
@@ C.2 AI For Science
The experimental results of using ResNet for image classification and ESM2 for protein sequence prediction demonstrate that the straightforward method proposed based on the AR framework is still generally effective in fields beyond NLP. However, it is important to note that in non-language modeling problems, the performance of the AR law does not exhibit the same strictness as it does in language modeling problems. While the overall trend aligns with the AR law, small fluctuations may occur locally. This is because the AR framework is based on relative overfitting for correction, which is quite intuitive in basic language modeling problems, but exhibits more complexity in other scenarios. Therefore, this paper posits that further in-depth research on relative overfitting and the AR law is necessary.
@@ D Complementary Experiments
In Appendix D, we demonstrate the performance of Pythia, RWKV, Mamba, and Qwen (based on int4 quantization) within the Qwen framework, which is based on the AR framework, on the fundamental language modeling benchmarks WikiText2, WikiText103, text8, and Pile (with Pile being a randomly sampled portion of the test set), as well as the 1BW benchmark. We found that the improvements under our framework are generally effective and largely conform to the AR law we proposed.
Qwen is a series of large language models developed by Alibaba Cloud in 2023. As a prominent commercial large language model, it competes with other industry models like GPT and Claude. The model architecture is based on the Transformer decoder framework, with optimizations for context length and instruction-following capabilities. We used five different sizes of the Qwen model: 0.5b, 1.5b, 3b, 7b, and 14b. To facilitate formatting, we will present the quantization results based on int4 in Table 8.
@@
Receptance Weighted Key Value (RWKV) is a novel architecture that combines the parallelizable training of Transformers with the constant-memory, linear-time inference of RNNs by leveraging a linear attention mechanism, achieving performance on par with similarly sized Transformers. In our experiments, we used the 169M, 430M, and 1.5B variants of RWKV.
@@
Mamba is a linear-time sequence modeling architecture that replaces traditional attention modules with input-conditioned Selective State Space Model (SSM) layers, enabling content-based reasoning over long sequences in O(n) time while maintaining Transformer-comparable accuracy. Mamba has demonstrated representative performance across modalities such as language, audio, and genomics. In our experiments, we used the 130M, 370M, 790M, 1.4B, and 2.8B variants of Mamba.
@@
Pythia is an open-source suite of autoregressive Transformer models developed by EleutherAI to enable rigorous scientific research into large language model behavior and training dynamics, with fully public data provenance and checkpointing. It is a highly classic series of scaling law models. In our experiments, we used the 14M, 70M, 160M, 410M, 1B, 1.4B, and 2.8B variants.
@@
A series of experimental results based on models with different structures and purposes indicates that our proposed AR framework, which is based on relative overfitting, has achieved widespread effectiveness, and the results are in almost complete accordance with the AR law we proposed after thousands of experiments.

### v4

@@ B.1 Larger models have more optional tokens
The number of tokens that the model may generate across all datasets increases with the increase in the model parameters.
B.2 Variation of Optimal Weight Ratio $\alpha$ with Parameter Ratio $k$
@@
Figure 4 shows that under the condition of the same main parameters $\theta_{LLM}$, the optimal weight ratio $\alpha$ for each model shows an increasing trend with increasing parameter ratio.
B.3 AR framework based on cosine similarity
As discussed in the main text, the AR framework emphasizes measuring judgment differences between models and ensembling relative overfitting information. Here, we present an alternative approach that, although not entirely successful, provides valuable insights into relative overfitting and the AR law.
We extract the embedding spaces from two models and compute the cosine similarity between each pair of tokens, storing these values for efficient retrieval during inference (see Algorithm1). We prioritize the LLM’s output during inference while using the SLM as a reference. Specifically, we obtain the probability distribution from the LLM and identify the token with the highest probability from the SLM. We then retrieve the cosine similarity vector between each token in the LLM’s vocabulary and the SLM’s highest probability token. Since these similarity values do not form a strict probability space as in our primary method, we employ element-wise multiplication to combine this similarity vector with the LLM’s output probabilities. To incorporate the relative overfitting information, we introduce a hyperparameter $p$ as the power exponent for the similarity vector (Algorithm2).
Algorithm 1 Computing Cosine Similarity Matrix
1:model, chunkSize
2:packedData, metadata
3$\triangleright$ $E$: Embedding matrix, $V$: Vocabulary size, $S$: Similarity matrix
4:$E\leftarrow\text{ExtractEmbeddings}(model)$
5:$E_{norm}\leftarrow\frac{E}{\|E\|_{2}}$triangleright Normalize embeddings
6:$V\leftarrow E_{norm}.\text{shape}[0]$
7:$S\leftarrow\text{zeros}(V,V)$
8:Compute similarity matrix by iterating through chunks:
9:for $i\leftarrow 0$ to $V-1$ step chunkSize do
10:     for $j\leftarrow i$ to $V-1$ step chunkSize do
11:         $S_{chunk}\leftarrow E_{i}\cdot E_{j}^{T}$
12:         if $i=j$ then
13:              $S[i:i+chunkSize,i:i+chunkSize]\leftarrow\text{upperTriangular}(S_{chunk})$
14:         else
15:              $S[i:i+chunkSize,j:j+chunkSize]\leftarrow S_{chunk}$
16:         end if
17:     end for
18:end for
19:$S_{quant}\leftarrow\lfloor S\cdot 15\rfloor$triangleright Quantize to 4-bit representation
20:return PackUpperTriangular$(S_{quant})$triangleright Store in compressed format
Algorithm 2 Similarity-Based Model Fusion
$model_{1}$, $model_{2}$, input, packedData, metadata, $p$ (power law hyperparameter)
2:Ensemble prediction probabilities
3$\triangleright$ $L$: Logits, $P$: Probabilities, $T$: Token indices, $S$: Similarity matrix
4:$S\leftarrow\text{UnpackUpperTriangular}(packedData,metadata)$triangleright Decompress similarity matrix
5:$L_{1}\leftarrow model_{1}(input)$
6:$L_{2}\leftarrow model_{2}(input)$triangleright Obtain logits from both models
7:$flat\_L_{1}\leftarrow\text{reshape}(L_{1},[-1,V])$
8:$flat\_L_{2}\leftarrow\text{reshape}(L_{2},[-1,V])$
9:$T_{2}\leftarrow\arg\max(flat\_L_{2},\text{dim}=1)$triangleright Identify model 2’s predicted tokens
10:$simWeights\leftarrow S[:,T_{2}]^{p}$triangleright Calculate similarity weights with power scaling
11:$P_{1}\leftarrow\text{softmax}(flat\_L_{1},\text{dim}=1)$
12:$P_{ensemble}\leftarrow P_{1}\odot simWeights$triangleright Fuse probabilities using similarity weights
13:$P_{ensemble}\leftarrow\frac{P_{ensemble}}{\sum P_{ensemble}}$triangleright Re-normalize to ensure valid probabilities
14:return $P_{ensemble}$
This approach is considered less successful because, despite implementing various optimization techniques, the storage requirements for the similarity matrix remain substantial without yielding performance improvements—in fact, it slightly underperforms compared to the method described in the main text. Furthermore, this approach still necessitates searching for appropriate hyperparameters under the AR law to ensemble the relative overfitting information between models effectively.
The primary value of discussing this method lies in exploring the AR law within the AR framework. Our experiments confirm that this approach still adheres to the pattern we proposed. Specifically, there exists an optimal value $p_{0}$ such that when $p<p_{0}$, performance gains increase with increasing $p$, while when $p>p_{0}$, performance gains decrease with increasing $p$. This observation further validates our theoretical framework.
The limited effectiveness of this more complex approach may be attributed to the fact that using similarity metrics and power-law adjustments as measures of judgment differences between models introduces additional complexity. Using only the highest probability token from the SLM may be insufficient to fully capture the nuanced relationships between models. This approach could achieve better performance improvements with more research and more sophisticated designs.
During our experimental process, we discovered that without introducing performance-related overfitting difference information between two models based on the AR framework, specifically the power-law hyperparameter p, we cannot guarantee that performance will consistently improve rather than decline. This further confirms our conclusion that the AR law, based on the AR framework and relative overfitting concept, is a prerequisite to ensure universal improvement rather than deterioration.
@@ C Experiments in the rest of the machine learning field
In this appendix, we detail the experimental setups employed. For image recognition, a suite of pre-trained ResNet models of varying sizes (ResNet-18 to ResNet-152) was utilized. Subsequently, these models were fine-tuned for one epoch in the CIFAR-10, CIFAR-10067 , and Tiny ImageNet-200 datasets 59 ; 60 ; 61 ; 62 ; 63 to assess classification efficacy across varying complexity levels. Currently, the ESM-2 model was used for the protein sequence prediction task. This task focuses on predicting protein sequences, which holds significant societal importance, including applications such as the design of novel protein sequences for therapeutic or industrial purposes. The predictive capabilities of the model were benchmarked in the UniRef5066 data set.
The experimental results demonstrate that our proposed ideas and framework can be universally applied to other areas of machine learning.
@@ C.1 Computer Vision
Our experiments employed a consistent post-training approach across CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. We fine-tuned ResNet series models (from ResNet18 to ResNet152) pre-trained on ImageNet. The fine-tuning process began by replacing the final fully connected layer, adjusting the output dimensions to match the respective datasets: 10 classes for CIFAR-10, 100 for CIFAR-100, and 200 for Tiny-ImageNet. All images were resized to 224×224 pixels and normalized using ImageNet parameters (mean=[0.485, 0.456, 0.406], std=[0.229, 0.224, 0.225]). We utilized SGD optimizer with a learning rate 0.001 and momentum of 0.9, typically running for one epoch. Performance was evaluated using Top-1 accuracy metrics for both individual models and ensembles.
@@ C.2 AI For Science
The experimental results of using ResNet for image classification and ESM2 for protein sequence prediction demonstrate that the straightforward method proposed based on the AR framework is still generally effective in fields beyond NLP. However, it is important to note that in nonlanguage modeling problems, the performance of the AR law does not exhibit the same strictness as it does in language modeling problems. Although the overall trend aligns with the AR law, small fluctuations may occur locally. This is because the AR framework is based on relative overfitting for correction, which is intuitive in fundamental language modeling problems, but exhibits more complexity in other scenarios. Therefore, this paper posits that further in-depth research on relative overfitting and the AR law is necessary.
@@ D Complementary Experiments
D.1 Supplementary Experiments of Chapter 4.1
In Appendix D, we demonstrate the performance of Pythia, RWKV, Mamba, and Qwen (based on int4 quantization) within the AR framework, which is based on the AR framework, on the fundamental language modeling benchmarks WikiText2, WikiText103, text8, and Pile (with Pile being a randomly sampled portion of the test set), as well as the 1BW benchmark. The improvements under our framework are generally practical and largely conform to the AR law we proposed.
Qwen is a series of large language models developed by Alibaba Cloud in 2023. As a prominent commercial large language model, it competes with other industry models like GPT and Claude. The model architecture is based on the Transformer decoder framework, with optimizations for context length and instruction-following capabilities. We used five different sizes of the Qwen model: 0.5b, 1.5b, 3b, 7b, and 14b. We will present the quantization results based on int4 in Table 8 to facilitate formatting.
@@
Receptance Weighted Key Value (RWKV) is a novel architecture that combines the parallelizable training of Transformers with the constant-memory, linear-time inference of RNNs by leveraging a linear attention mechanism. It achieves performance on par with similarly sized Transformers. We used 169M, 430M, and 1.5B of RWKV in our experiments.
@@
Mamba is a linear-time sequence modeling architecture that replaces traditional attention modules with input-conditioned selective state-space model (SSM) layers, enabling content-based reasoning over long sequences in O(n) time while maintaining Transformer-comparable accuracy. Mamba has demonstrated representative performance across modalities such as language, audio, and genomics. In our experiments, we used the 130M, 370M, 790M, 1.4B, and 2.8B variants of Mamba.
@@
Pythia is an open-source suite of autoregressive Transformer models developed by EleutherAI to enable rigorous scientific research into large language model behavior and training dynamics, with fully public data provenance and checkpointing. It is a highly classic series of scaling law models. We used the 14M, 70M, 160M, 410M, 1B, 1.4B, and 2.8B variants in our experiments.
@@
A series of experimental results based on models with different structures and purposes indicates that our proposed AR framework, which is based on relative overfitting, has achieved widespread effectiveness, and the results are in almost complete accordance with the AR law we proposed after tens of thousands of experiments.
D.2 Supplementary Experiments of Chapter 4.2
The utilization of models trained on diverse datasets within the main body of this work is primarily intended to offer a more granular illustration of the relative overfitting phenomenon under various conditions. This section introduces an approach for Large Language Model (LLM) rectification employing a singular auxiliary model. Specifically, we trained and fine-tuned a model across all datasets utilized in our experiments, designed to exhibit performance slightly inferior to the primary model on most benchmarks. This auxiliary model shares the same architecture (a simple Transformer) and parameter count (approximately 180M) as the primary model. Notably, we intentionally engineered its performance on the 1BW dataset to be marginally superior to that of the main model. The corresponding results are presented in Table 10. It is pertinent to mention that the performance achieved on the 1BW dataset is still considerably below its potential optimum. Naturally, such non-generically trained models, through iterative fine-tuning, may undergo complex and multifaceted changes. Consequently, relative overfitting in these instances can be more intricate than models trained systematically according to established scaling laws. A more profound investigation into these complex scenarios remains a subject for future research.
Model WikiText2 WikiText103 LAMBADA text8 1BW
GPT2-Train 15.24 17.74 37.63 1.12 44.21
Other-Train 12.85 23.97 32.14 1.04 22.79
GPT2-1.5b 55.26% 54.44% 73.96% 25.43% 63.84%
Pythia-2.8b 58.92% 53.51% 64.84% 33.15% 71.89%
Mamba-2.8b 57.13% 51.52% 63.29% 31.95% 69.91%
RWKV-1.5b 62.25% 57.45% 66.26% 34.18% 72.98%
D.3 Supplementary Experiments of Chapter 4.5
Here, we will present more results regarding the AR law. Experimental results indicate that our integration theory is universally valid. In fact, in all of our experiments, with the exception of a very few occurrences of one in ten thousand, the AR law holds universally.
Figure 6: The AR law of the Pythia model.
Figure 7: Comparison results of the AR law for four different architectural combination models.
Figure 8: The AR law of the Pythia model.